Solve by completing the square:
y^2-6y-8=0
Your help is truly appreciated! Thanks!

Answer by ankor@dixie-net.com(22740)

Solve by completing the square:
y^2 - 6y - 8 = 0
:
y^2 - 6y + ___ = + 8
:
Complete the square;
y^2 - 6y + 9 = 8 + 9
:
(y - 3)^2 = 17
:
Find the square root of both sides:
y - 3 = +/-Sqrt(17)
:
y = 3 + Sqrt(17)
and
y = 3 - Sqrt(17)
:
Did this help, any questions